OpenStudy (gabylovesyou):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

b^2-4ac. If b^2-4ac is a perfect square then it implies that roots are rational and hence u can factor up the equation. Comparing with ax^2+bx+c = 0, a = 2, b = 7, c = 3 b^2-4ac = 7^2 - 4.2.3. = 49-24 = 25.